OpenStudy (anonymous):

Find all values of x for which the function is Differentiable at P(x) = sin (| x |) -1

OpenStudy (anonymous):

i doubt that

OpenStudy (anonymous):

if \(x<0\) then \(\sin(|x|)=\sin(-x)\) and so the derivative would be \(-\cos(x)\)

OpenStudy (anonymous):

\[D_x \sin(x)=\cos(x)\\ \begin{cases}|x|=x, \;\;x\geq0 \\ |x|=-x, \;\;x<0 \end{cases}\\ \cos(-x)=cos(x)\\ \sin(-x)=-\sin(x) \] If x < 0 Dx[sin(|x|)]=Dx[sin(-x)]=-cos(x)

OpenStudy (anonymous):

and since at 0 \(-\cos(0)\neq \cos(0)\) it is not differentiable there

OpenStudy (anonymous):

it is not continuity it is differentiability